OpenStudy (anonymous):

Which sequence is geometric and has 14 as its fifth term and 12 as the common ratio?

OpenStudy (welshfella):

the 5th term = ar^4 where a = first term and r = common ratio so 14 = a * 12^4

OpenStudy (welshfella):

14 = 20736 * a a = 14/20736

OpenStudy (anonymous):

…, 1, 12, 14, 18, … …, 14, 12, 1, 2,… …, 172, 136,118, 19, … …, 8, 4, 2, 1, …

OpenStudy (welshfella):

its neither of these are you sure you have the question right?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

something's wrong somewhere